geerky42 (geerky42):

How many values of \(\theta\) satisfy \(2\cos(3\theta-1) = 0\), where \(-\pi \le \theta \le \pi\)?

Parth (parthkohli):

Divide both sides by \(2\), and you get\[\cos(3\theta - 1) = 0\]For what all angles is cosine zero?

geerky42 (geerky42):

I got 2 but the answer is 6. I don't understand...

Parth (parthkohli):

Consider the graph of \(\cos(x)\).

geerky42 (geerky42):

\(\dfrac{\pi}{2}\) and \(-\dfrac{\pi}{2}\)

Parth (parthkohli):

You're right! o.O

geerky42 (geerky42):

so 3θ−1 got to be equal to either \(\dfrac{\pi}{2}\) or \(-\dfrac{\pi}{2}\) so i found 2 values theta can be, but key answer said 6...

Parth (parthkohli):

\[\cos(3\theta - 1) = \cos(1 - 3\theta) = 0\]

Parth (parthkohli):

Now, do you get it? ;-)

OpenStudy (mertsj):

I you want all the theta answers between pi and -pi, you must find all the 3theta answers between 3pi and -3pi

geerky42 (geerky42):

Wait, I dont get it...

Parth (parthkohli):

What Mertsj said is what I wanted to eventually point out...

Parth (parthkohli):

I don't want to gobble this for you, but there's this thing called "frequency", which decreases when you increase the coefficient of \(\theta\).

Parth (parthkohli):

So if the graph of cosine is likeIt'd becomeMore close to each other.